A polyhedron has 7 faces and 10 vertices. How many edges does the polyhedron have?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

For any polyhedron,

F + V – E = 2

Given here, F = 7, V = 10, E = ?

Substituting the values, we get;

7 + 10 – E = 2

17 – E = 2

E = 17 – 2

E = 15


Thanks